### v1

@@ B.4 Implementation Details
We simulate parallel training by time multiplexing and using preallocated replay and communication tensors. The experiments were run on a variety of hardware, including mobile CPUs (Apple M2) and workstation GPUs (on a single NVIDIA RTX A5000 and single-GPU, and on a single NVIDIA Tesla K40).
Appendix C Choice of Baseline
C.1 SplitNN as a Baseline
In this work we present DVFL, an alternate strategy to VFL. We demonstrate its prowess through a DVFL implementation for ANNs. Since DVFL is altogether a new strategy to train on vertically partitioned data, it makes sense to benchmark it against the current status quo for VFL with ANNs, which is SplitNN Gupta & Raskar (2018); Ceballos et al. (2020).
SplitNN with a timeout
In its base form, the SplitNN algorithm does not have any explicit fault handling strategy. Implicitly (assuming that the coordination between hosts and guests is interrupt-driven), the hosts use the “wait” strategy. Training stops completely if inputs are missing. We also consider strategies using “timeout”. Timeout allows training to continue, although hampered. We handle timeouts by skipping or replacing the values. If a message appears after the deadline, it needs to be discarded lest it appear as if it were intended for the next iteration. This artificial baseline represents a best-case fault-tolerance for vanilla VFL. This way, it gets a fair shot against DVFL. Indeed, since DVFL suffers from some information loss (since feedback is very limited), VFL equipped with these strategies may still outperform DVFL at low fault rates.
Busy-waiting SplitNN
Under “wait”, training can be made to restart if the failed participant comes back online by busy-waiting viz. the repeatedly polling a failed guest. If the guests send sample IDs along with their features (and hosts cache messages until they are used), they may figure out which iteration a particular message is intended for. This way, no information is lost. However, until that failed participant is back online, training is held up. This is also a poor strategy from a systems perspective, since it wastes bandwidth and computational resources, especially when there is a message authentication in place. If a guest is laggy and the host polls it many times before it repsonds, the cache of messages the host must store will be very large. Neither in this case nor the timeout case is there a well-defined strategy to handle backward-pass faults.
C.2 Other Possible Baselines
DVFL addresses various VFL issues such as fault tolerance, privacy and sample size with one fell swoop. There are various mechanisms built upon standard VFL algorithms (including SplitNN) which aim to address some or the other of these issues. FedVS Li et al. (2023) is a work that addresses multiple, i.e. fault tolerance and privacy. FedVS might be a reasonable benchmark for DVFL, and we have presented some results against its paper in Appendix A. DVFL outperforms FedVS in perfect conditions. In the main paper, we showed that DVFL performance does not degrade much with faults with our generalized fault model. On the other hand, FedVS performance is completely fault tolerant, except under a much more stringent fault model. Beyond this, FedVS constrains guest models to use a specific type of model– polynomial networks Livni et al. (2014a; b). Moreover, FedVS only deals with faults during the forward pass. FedVS considers a restricted fault model: upto a fixed number of guests can be late to/fail to communicate their activations with the hosts. These are also true for most of its predecessors e.g. Chen et al. (2020a); Hu et al. (2019); Shi et al. (2022). DVFL does not require any such system parameter, and is tolerant to a broader class of faults where any participant can fail at any time, including during backpropagation.
Appendix D Communication and Space Complexity
D.1 Periodic Communication
@@
The owner initiates every round of training as well as communication for all agents. Since the training of guests and hosts are decoupled, the owner may ask guests to communicate their latent activations (and therefore start host training rounds) at arbitrary intervals. In VFL, guest model updates are dependent on host feedback. So there is every-shot communication, i.e. guests send every single output to the host.
@@
With one-shot communication such as FedOnce Wu et al. (2022), the volume of communication is greatly reduced, at the cost of model performance. Equivalently, Belilovsky et al. (2020); Siddiqui et al. (2023) demonstrate that updating all layers on local objectives on each forward pass performs better than fully training and freezing a layer before moving on to the next.
@@
Note that the dimension of the latent representations communicated by each guest to reach host is $80$, and each dimension corresponds to a 32-bit floating point number. The MNIST training dataset has $60000$ samples. If the total number of guest training epochs is $N$ and the period of communication epochs is $K$, the total communication cost per guest is ||×$|\mathcal{H}|\times\lfloor\frac{N}{K}\rfloor\times 60000\times 80\times 32$ bits. Each message received is saved by the host in the activation replay [^3] Although this is not necessary if the host will not reuse activations, viz. a host wants to train for $k$ iterations and it receives activations of $k$ iterations from the guests. This is discussed in more detail in Appendix D.2.
0055551010101015151515202020202525252596.996.996.996.99797979797.197.197.197.197.297.297.297.297.397.397.397.397.497.497.497.4Number of HostsMNIST Test Accuracy (%)
@@
D.2 Activation Replay Mechanism & its Computational Overhead
A host reads its model inputs from its activation replay buffer. Thanks to this, it can train asynchronously to the guests. The host can train independently to the guests as long as the guests have communicated at some prior point of time. In fact, that is the case in our experiments– we simulate host models’ training offline, after guest training. As discussed in Appendix D.1, this allows control over when the guests communicate.
Another way to look at this is: hosts do not need to train for the same number of iterations as guests communicate messages for. This includes the case where guests communicate every iteration, but hosts want to train more (or less) iterations. As we detailed in Appendix B.3, we actually implement this for our experiments. While we use 40 epochs to train the entire SplitNN model, we only use 20 epochs to train the DVFL guests. Guests need not remain online for any more time than required to train 20 epochs. So, not only does this allow for more flexibility, it also helps fault tolerance.

### v2

@@ D.4 Implementation Details
We simulate parallel training by time multiplexing and using preallocated replay and communication tensors. The experiments were run on a variety of hardware, including mobile CPUs (Apple Silicon M2) and workstation GPUs (on a single NVIDIA RTX A5000 and on a single NVIDIA Tesla K40).
Appendix E More Discussion on Related Works
E.1 Fault-Tolerant Distributed Optimization
The impact of redundancy for fault tolerance in distributed systems is well studied. Liu et al. (2021) provide a rigorous analysis in the case of distributed optimization, showing that systems which are $2f$-redundant are approximately resilient to $f$ Byzantine agents. The authors also present two fault-tolerant algorithms for horizontally distributed convex optimization. Liu (2021); Bouhata et al. (2022) extensively review approaches to fault tolerance in distributed learning. Most methods require synchronization by an aggregator and incur computational overhead. Further, most methods involve manipulating gradients, which forces a choice between convergence rate and tightness.
For VFL: A limited number of works exist to tackle the related problem of straggling participants.
Li et al. (2023) introduce FedVS, an algorithm for systems where guest models are polynomial networks (Livni et al., 2014a; b). FedVS uses a form of secret sharing to introduce redundancy in the input data. The redundancy is so arranged such that, when there are at most a fixed number of failed or late guests, timely guests have enough combined information to account for the missing messages.
In situations where multiple participants might have labels (i.e. there is a horizontal federation of label owners), (Xia et al., 2023) exploits redundancy and periodic synchronization to provide straggler resistance. One specially designated horizontally federated host is responsible for orchestrating the synchronization between hosts. Typically, straggler-resistant algorithms are either designed to tolerate a fixed number of stragglers or a finite duration of lag (i.e. a crash would still be catastrophic). Further, FedVS and its predecessors (Shi et al., 2022; Chen et al., 2020a; Hu et al., 2019) typically only address faults in the forward pass.
E.2 Federated Transfer Learning
Federated Transfer Learning (FTL) is a strategy for systems where there is some overlap in both sample space and feature space. The host owns labels and some features (i.e. it is an active participant) for entities in its sample space. Guests have limited labels for their task, but have some small overlap in feature space and sample space with the host (“aligned” samples).
Liu et al. (2020) present Secure Transfer Federated Learning (STFL), a framework for a single host and a single guest. Both parties learn a common latent space using local feature extractors. Encodings of the passive participant’s unaligned samples in this feature space is used by the passive participant to learn its task model. Feng & Yu (2020); Feng et al. (2022) extend this idea to multiple parties.
Much like DVFL, these methods exploit an unsupervised latent space to learn from data that is not present in the intersection. However, unlike DVFL, these methods exploit the existence of common feature space between participants. This problem is relevant to robust distributed sensor fusion (Liu, 2021), where there is some redundancy/overlap in the “view” of each participant.
For the related case of VFL with small $|\mathcal{S}_{\mathrm{guests}}\cap\mathcal{S}_{\mathrm{labels}}|$, Sun et al. (2023) introduce few-shot Vertical Federated Learning, where a few epochs of training on the aligned dataset are followed by semi-supervised learning on the unlabelled entities of each client’s dataset. To generate labels, guests cluster the gradients they receive from the aligned samples. While this enables remarkable model performance, there is unwanted computational overhead required to generate the labels. Moreover, this is a form of label inference attack! Not only is this a breach of privacy, this clustering could be used downstream by a malicious agent to perform feature inference attacks (see §2.4.3).
E.3 Greedy and Localized Learning
Algorithms which update parameters in hidden layers of NNs based on local or auxiliary objectives (as opposed to the global label loss) have recently seen renewed interest. In the scope of VFL training, they enable some useful properties for NN training.
By eschewing the wait to evaluate the global objective, most greedy algorithms are backward unlocked if not update unlocked. Belilovsky et al. (2020) propose a replay buffer mechanism to partially address forward locking for model parallel training. Each agent responsible for training a module uses a replay buffer to store activations from the previous module’s agent. During a forward pass, the agent reads an activation from the buffer. If the forward signal has propagated through the previous module, it would have written the latest activation to the buffer. If not, the previous activation is reused.
Appendix F Communication and Space Complexity
F.1 Periodic Communication
@@
The owner initiates each round of training as well as communication for all agents. Since the training of guests and hosts are decoupled, the owner may ask guests to communicate their latent activations (and therefore start host training rounds) at arbitrary intervals. In VFL, guest model updates are dependent on host feedback. So there is every-shot communication: guests send every single output to the host.
@@
With one-shot communication such as FedOnce (Wu et al., 2022), the volume of communication is greatly reduced, at the cost of model performance. Equivalently, Belilovsky et al. (2020); Siddiqui et al. (2023) demonstrate that updating all layers on local objectives on each forward pass performs better than fully training and freezing a layer before moving on to the next.
@@
Note that the dimension of the latent representations communicated by each guest to reach host is $80$, and each dimension corresponds to a 32-bit floating point number. The MNIST training dataset has $60000$ samples. If the total number of guest training epochs is $N$ and the period of communication epochs is $K$, the total communication cost per guest is ||×$|\mathcal{H}|\times\lfloor\frac{N}{K}\rfloor\times 60000\times 80\times 32$ bits. Each message received is saved by the host in the activation replay [^5] Although this is not necessary if the host will not reuse activations, viz. a host wants to train for $k$ iterations and it receives activations of $k$ iterations from the guests. This is discussed in more detail in the Appendix F.2.
005555101010101515151520202020252525259797979797.297.297.297.297.497.497.497.497.697.697.697.697.897.897.897.8Guest training epochs per communication epochMNIST Test Accuracy (%)
@@
F.2 Activation Replay Mechanism & its Computational Overhead
A host gets its input from its activation replay buffer. That allows it to train asynchronously wrt the guests. The host can train independently of the guests as long as the guests have communicated earlier. Indeed, that is the case in our experiments– we simulate host models’ training offline, after guest training. As discussed in Appendix F.1, this allows control over when the guests communicate.
Another way to look at this is: hosts do not need to train for the same number of iterations as guests communicate messages for. This includes the case where guests communicate every iteration, but hosts want to train more (or less) iterations. As we detailed in Appendix D.3, we actually implement this for our experiments. While we use 40 epochs to train the entire SplitNN model, we only use 20 epochs to train the DVFL guests. Guests need not remain online for any more time than required to train 20 epochs. So, not only does this allow for more flexibility, it also helps fault tolerance.
@@ F.2 Activation Replay Mechanism & its Computational Overhead
F.3 Empirical Time Complexity
We compare the following regimes:
  1. 1.
    DVFL
    1. (a)
      MSE loss
      • Same setup as main experiments
      • Guests trained for 20 epochs
      • Hosts trained for 40 epochs
      • Owner trained for 60 epochs
    2. (b)
      SimCLR loss
      • Guests trained for 5 epochs
      • Owner trained for 60 epochs
      • Guests optimized using Adam with $4\times 10^{-3}$ learning rate, $\beta_{1}=0.9$, $\beta_{2}=0.999$, weight decay $10^{-5}$
      • $\text{linear}(160,50)\to\text{batchnorm}\to\text{ReLU}\to\text{linear}(50,50)\to\text{batchnorm}\to\text{ReLU}\to\text{linear}(50,50)$ projector
      • No host
      • Owner model is equivalent to the host model and owner model of our main experiments in sequence
      • Owner is optimized using SGD with learning rate $10^{-1}$, momentum 0.9
      • Batch size of 600.
  2. 2.
    VFL
    1. (a)
      No localized loss
      • Same setup as main experiments
      • System trained for 60 epochs
We measure the CPU process time on an Apple Silicon M2 CPU and compare the following process + system times:
  • Forward: time per epoch spent in the forward pass (including loss calculation, if any)
  • Backward: time per epoch spent in the backward pass (including optimizer step)
  • Wait: time per epoch spent waiting for gradients
  • Process: total time the process must be alive
We do not include communication delay or spent managing tensors, as those are heavily implementation-dependent. We also do not count data augmentation time for SimCLR, since that may be precomputed before guest-host training.
Table 8: DVFL allows guests to train without wait and terminate their processes much sooner than VFL. Note that SimCLR requires a large batch size (here: 600), and therefore gets swapped out of memory frequently during training. The time reported includes that.
Strategy
DVFL (MSE) DVFL (SimCLR) SplitNN
Quantity Guest forward (s/ep) 2.144 4.872 1.830
Guest waiting (s/ep) 0.000 0 2.180
Guest backward (s/ep) 0.650 10.488 0.001
Host forward (s/ep) 1.748 4.657 0.982
Host backward (s/ep) 0.806 20.778 0.274
Owner forward (s/ep) 1.271 1.149
Owner backward (s/ep) 0.564 2.430
Guest process (s) 55.883 305.596 80.220
The accuracies observed are:
  • SplitNN 97.85%
  • DVFL (MSE) 97.80%
  • DVFL (SimCLR) 97.78%
Accuracies on SimCLR under faults (0 hosts):
  • Guest
    1. 1.
      Rejoin Rate = 1.0: 97.63%
    2. 2.
      Rejoin Rate = 0.5: 97.64%
    3. 3.
      Rejoin Rate = 0.1: 97.41%
  • Communication
    1. 1.
      Rejoin Rate = 1.0: 97.71%
    2. 2.
      Rejoin Rate = 0.5: 97.63%
    3. 3.
      Rejoin Rate = 0.1: 97.52%
The batch size is larger, so losing a single forward pass loses more information.